OpenStudy (anonymous):

A sign is in the shape of a rhombus with a 60° angle and sides of 9 cm long. Find its area to the nearest tenth.

OpenStudy (whpalmer4):

Formula for the area of a rhombus with angle \(\theta\) and side length \(a\) is \[A = 2a^2\sin \theta \cos \theta\]